OpenStudy (anonymous):

Find the domain of each function f(x)=1-e^(x^2)/1-e^(1-x^2)

OpenStudy (anonymous):

is your denominator \[1-e^{1-x^2}\]?

OpenStudy (anonymous):

if so, set \[1-e^{1-x^2}=0\] and solve for \(x\) you get \[e^{1-x^2}=1\] making \[1-x^2=0\] and so \[(1+x)(1-x)=0\] solve for \(x\) and say all real numbers except those

OpenStudy (anonymous):

hope the second step was clear, only way \(e^{\text{something}}=1\) is if \(\text{something}=0\)

OpenStudy (anonymous):

yes it is.

OpenStudy (anonymous):

ok good, so you have the answer right? and the method is ok?

OpenStudy (anonymous):

How would I put it in interval notation.. (-inf,-1)U(1,inf)?

OpenStudy (anonymous):

oh no you missed the part in between

OpenStudy (anonymous):

:/ uh oh lol

OpenStudy (anonymous):

\[(-\infty, -1)\cup (-1,1)\cup (1,\infty)\] which is a pretty damn long winded way of saying everything except -1 and 1

OpenStudy (anonymous):

I was close, okay thanks! :)